Miriam Co has a bu=1, Wd=30%, T= 21% RPM= 5%, rRF=3%. Using the CAPM model, calculate Miriam Co's cost of equity?
Answer #1

Beta unlevered =1
Beta Levered =Beta Unlevered *((1+(1-Tax Rate)*Debt/Equity)=1*((1+(1-21%)*30%/70%) =1.3386
Cost of equity =Risk free rate+Beta*RPM =3%+1.3386*5% =9.69%